OpenStudy (anonymous):

how to get $$\int_0^\infty \ x^n d\frac{(e^{-ax})}{-a}$$ from $$\int_0^\infty \ x^n {(e^{-ax})} {dx}$$

OpenStudy (anonymous):

Thank you

OpenStudy (irishboy123):

do the sub: \[u = -\frac{e^{−ax}}{a}; \ du = -\frac{1}{a}(-a e^{-ax}) \ dx; \ dx = \frac{ du}{e^{-ax} }\] and switch the limits round

OpenStudy (anonymous):

Thank you!!. It is based on reduction formula